solve by using any method
5(x+y)=x-4
-y= 4x+4/5

Answer by stanbon(57424)

solve by using any method
5(x+y)=x-4
-y= (4x+4)/5
-----------------
Rearrange:
4x + 5y = -4
4x + 5y = -4
------
Subtract 1st from 2nd
to get
0 = 0
----
Solution: All points of the form (x, -(4/5)(x+1)
---------------------------
Cheers,
Stan H.